OpenStudy (anonymous):

The number of books checked out from the library varies directly with the number of book reports assigned. There were 3375 books checked out when there were 225 book reports assigned. How many book reports have been assigned if there are 13,500 books checked out of the library?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

it's direct correlation. 3375 books checked out -> 225 book reports have been assigned 13500 books checked out -> x book reports have been assigned so; 3375/13500=225/x 1/4=225/x x=900 do you understand??

OpenStudy (anonymous):

yes thank you(:

OpenStudy (anonymous):

welcome :)